OpenStudy (anonymous):

Which equation represents the line passing through the points (3, 2) and (−9, 6)? x − 3y = 9 x + 3y = 9 3x − y = −9 3x + y = 9

OpenStudy (anonymous):

we can use slope intercept form

OpenStudy (anonymous):

y=mx+b?

OpenStudy (anonymous):

no (y-y0)=m(x-x0)

OpenStudy (anonymous):

here (x0,y0) is one of those points and m is slope

OpenStudy (anonymous):

slope=y2-y1/x2-x1

OpenStudy (anonymous):

so slope will be 6-2/-9-3 =1/-3 =-1/3

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

by using slope intercept form,we have (y-6)=(-1/3)(x--9)

OpenStudy (anonymous):

are you getting it?

OpenStudy (anonymous):

mostly yes

OpenStudy (anonymous):

so we will now expand this

OpenStudy (anonymous):

y-6=(-1/3)(x+9) --->y-6=(-x/3) -9/3

OpenStudy (anonymous):

y+x/3 =-3+6 --->y+x/3=3

OpenStudy (anonymous):

any doubts till here?

OpenStudy (anonymous):

nope, but it's the 2nd one, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u multiply this by 3..and will get the answer as the second one

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

You are welcome:D